OpenStudy (anonymous):

Solve for X: 8^(3x)=16^(x-5)

OpenStudy (anonymous):

x = -4

OpenStudy (anonymous):

i need to know how u got that not just the answer.

OpenStudy (anonymous):

rewrite 8 and 16 as powers of 2

OpenStudy (anonymous):

8^(3x)=2^(9x) 16^(x-5)=2^(4(x-5)) Now solve 9x=4x-20 -> 5x=-20 -> x=-4

OpenStudy (ghazi):

(2)^3*3x=2^4*(x-5)...equate the powers ....9x= 8(x-5)

OpenStudy (ghazi):

since 2^3=8 and 2^4=16